The time period of a satellite of earth is 5 hours If the separation between the earth and the earth and the satellite is increased to 4 times the previous value, the new time period will become
(a) 10 hours
(b) 80 hours
(c) 40 hours
(d) 20 hours

Grade:6

3 Answers

Deepak Patra
9 years ago
(c) According to Kepler’s law of planetary motion T2 ? R3 ? T2 = T1 (R2/R)3/2 = 5 x [4R/R]3/2 = 5 x 23 = 40 hour
Chahat Gutpa
6 years ago
Time period = total distance/velocityT=2πr/√grT`=2π(4r)/√g(4r)T`=4(2πr)/2√grT`=2(2πr)/√grT`=2TT`=2×5T`=10 hours(option A)
Chahat Gutpa
6 years ago
Time period = total distance/velocity T =2πr/√gr T`=2π(4r)/√g(4r) T`=4(2πr)/2√gr. T`=2(2πr)/√gr T`=2T T`=2×5T`=10 hours(option A)